A projectile is thrown at 60° with the horizontal. After 4s it is moving perpendicular to its initial direction of motion. If g = 10 m/s^2 , then initial speed of throw is

Grade:11

1 Answers

Prashant Kumar
6 years ago
For this type of questions Simply use the formulaT=u/g*sinθSol. T=4θ= 60°Substitute the values4= u/10* sin60°Sin60°= √3/2=>U= 4*10*√3/2 = 20√3 m/sAns correct